Consider a die for which it is postulated that an odd face is twice as likely to occur as an even face. Based upon this assumption, together with the axioms of probability, find the probabilities of obtaining the following events on a single toss:

• {1}

• {2}

• {2, 4, 6}

• {1, 2, 3}
Answer #1

here as we know that sum of probability of events on sample space is 1

P(1)+P(2)+P(3)+P(4)+P(5)+P(6)=1

also P(1)=P(3)=P(5)=2*P(2)=2*P(4)=2*P(6)

solving above P(1)=P(3)=P(5)=2/9 and P(2)=P(4)=P(6)=1/9

hence below are asked probabilities:

1) P(1) =2/9

2) P(2)=1/9

3)P(2,4,6)=P(2)+P(4)+P(6)=(1/9)+(1/9)+(1/9)=3/9=1/3

4)P(1,2,3)=P(1)+P(2)+P(3)=(2/9)+(1/3)+(2/9)=5/9